Circular No. 5972

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. Telephone 617-495-7244/7440/7444 (for emergency use only) TWX 710-320-6842 ASTROGRAM CAM EASYLINK 62794505 MARSDEN@CFA or GREEN@CFA (.SPAN, .BITNET or .HARVARD.EDU)

SUPERNOVA 1994K IN ANONYMOUS GALAXY J. Mueller reports her discovery of a supernova of red mag ∼ 17.5 on

a plate taken with the 1.2-m Oschin Schmidt Telescope by D. Griffith and J. D. Mendenhall on Apr. 1 in the course of the second Palomar Sky Survey. The supernova is located at α = 9^{h}57^{m}43^{s}.09, δ = +13^{o}06′57′′.3 (equinox 1950.0), which is 22′′ west and 15′′ north of the galaxy’s nucleus.

B. Schmidt and R. Kirshner, Harvard-Smithsonian Center for Astrophysics (CfA), report that a spectrum (320–850 nm) taken by J. Huchra (CfA), C. Heller (University of Kentucky), and S. Muscarella (CfA) with the Multiple Mirror Telescope (+ Blue Channel) shows this object to be of type Ia, ∼ 6 weeks past maximum light. The redshift of the host galaxy, derived from its absorption lines, yields a velocity of recession of 7000 km/s.

SUPERNOVA 1994I IN NGC 5194 A. Clocchiatti, M. Brotherton, R. P. Harkness, and J. C. Wheeler,

University of Texas, report: “A fully-reduced spectrogram (range 459–998 nm, resolution 1.2 nm) of SN 1994I, obtained on Apr. 8.24 UT at the 2.7-m telescope of McDonald Observatory, shows several P-Cyg profiles superimposed on a blue continuum. Proposed identifications of the features and wavelengths of the blueshifted absorptions are: Fe II at 478.0 and 496.0 nm; He I 587.6-nm at 569.3 nm, perhaps blended with Na D; Si II 635.5-nm at 616.5 nm; O I 777.4-nm at 755.1 nm; and Ca II 849.8-, 854.2-, and 866.2nm at 817.8 nm. Superimposed on the redshifted maximum of the feature at 616.5 nm (Si II), there are two weak and broad absorption troughs with minima at 639.2 and 647.7 nm. If the feature at 647.7 nm is associated with He 667.8-nm, then its velocity matches that of He I 587.6-nm (∼ −10 000 km/s). The He features are weak and the spectrum broadly resembles the spectra of SN 1983V and SN 1987M at early phases. SN 1994I, then, should be classified as a type-Ic event (cf. IAUC 5966). Observations over all wavelengths are strongly recommended, especially of Paα and Paβ, lowresolution infrared spectroscopy targeting He I 108.3- and 206.0-nm, and high-resolution optical spectroscopy of the strong interstellar absorption Na D lines, which appear redshifted by ∼ 420 km/s.”

H. Mikuž, Ljubljana, Slovenia, provides the following CCD photometry (comparison-star magnitudes from H. Corwin): Apr. 7.88 UT, V = 13.05 ± 0.03, B − V = +0.51 ± 0.05.

1994 April 11 Daniel W. E. Green